6. Randomly a card is drawn from a deck of ordinary playing cards. You win LKR 1000 the card is a spade or an ace. What is the probability that you will win the game?

A. 1/13

B. 13/52

C. 4/13

D. 17/52

E. 18/52


1
Expert's answer
2021-10-04T18:25:23-0400
"P(S\cup A)=\dfrac{13+4-1}{52}=\dfrac{4}{13}"

C. 4/13